Fermi National Laboratory


Accelerator Update for December 7, 2005 - December 9, 2005

Wednesday December 7

The day shift began with the Tevatron (TeV) sector B1 separator bake out continuing, with the Antiproton Source (Pbar) conducting studies, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Cryo system technicians from CHL reported at 8:14 AM that the TeV sector A4 cryo system was ready to go.

At 8:16 AM, a Linac quadrupole power supply went out of tolerance. A Linac tech made adjustments and had it back up by 8:55 AM.

NuMI turned off for an hour at 9:36 AM while interlock techs conducted system checks.

Water techs began replacing the gasket for a CUB water column at 1:46 PM. This was holding off the TeV turn on due to no LCW.

A Main Injector (MI) expert began setting up for a MI/NuMI slipstacking study at 2:37 PM.

Operations successfully completed a TeV dump test and a QBS test at 3:02 PM.

At 3:35 PM, Operations reported that the TeV quenched on its first training ramp at 948 GeV.

Around 4 PM, the water group reported that they had replaced the gasket on column 3 at CUB, but that didn’t resolve the problem; more work was needed.

Operators interlocked the Collider Detector at Fermilab (CDF) at 5:26 PM.

The MI study ended at 6:28 PM.

At 6:41 PM, a TeV VCB popped open and wouldn’t reset. An expert came in to investigate.

At 9 PM, Operations reported that the TeV sector B2 VCB work was unsuccessful. More work will continue on Thursday.

Thursday December 8

The midnight shift began with TeV B1 separator cooling off from its bake out, with Pbar conducting studies, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations reported that they had ramped the TeV to 980 GeV at 12:43 AM, but a minute later the TeV suffered another sector B1 quench.

Operations successfully ramped the TeV at 5:05 AM.

Operations put the TeV through a dry squeeze at 5:34 AM. Operators found that the D-Zero (D0) dipoles were not ramping.

Operations put the TeV through another dry squeeze at 6:17 AM.

Operations prepared the Switchyard (SY), CDF, and D0 for access at 7:15 AM.

A MI expert began a four hour 2.5 Mhz study at 8:28 AM.

Experts reported that the TeV sector B2 VCB was fixed at 11:10 AM.

An RF expert reported at 4:18 PM that the TeV RF was up and running.

The TeV turn on had begun when operators reported at 6:44 PM that the TeV sector B2 VCB wouldn’t close. A TeV expert recommended bypassing the supply again. It will be checked on Friday.

At 6:53 PM, Operations reported that a SY power supply (HP3DS) wouldn’t turn on. Operators called in a power supply expert to investigate.

At 10:21 PM, Pbar experts reported having trouble getting beam into the Debuncher. D:IB had tripped off on a ground fault indication. Operations called in an expert who reported a power supply problem that would have to wait until the morning for repair.

Friday December 9

The midnight shift began with TeV turn on continuing, with Pbar off due to problems with D:IB, and with MiniBooNE and NuMI taking beam, and with MCenter off due to HP3DS problems.

The power supply expert investigating the HP3DS problem reported at 1:01 AM that the repairs would have to wait until the morning.

Operations put the TeV through a dry squeeze at 5:11 AM.

Operators began preparing the TeV for an access into sector B1 to remove equipment.

The Plans for Friday and the Weekend

The plans are to work on the TeV, Pbar, and Switchyard. Hopefully, the first store will be in the TeV by late Sunday.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/9/2005   email Fermilab